OpenStudy (anonymous):

2-3x = l5x-3I It's an absolute value equation.

OpenStudy (yttrium):

You can actually split this equation into two cases: 2-3x = 5x -3 and -(2x-3) = 2x - 3 There you can solve for the two values of x. :)

OpenStudy (john_es):

\[2−3x=5x−3 \text{ if } x>3/5\] and \[2x−3=−(5x−3) \text{ if } x<3/5\] Copy paste troubles.

OpenStudy (zpupster):

Case1: 5x-3 = 2-3x Case 2: 5x-3 = 3x-2 solve for these equations because 5x-3 could be positive or negative

OpenStudy (anonymous):

Thank you so much guys!